# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ A.4 Section 4
As $c\cap W_{A}=\bigcup_{\omega\in A}(c\cap W_{\omega})$, we infer from lemma A.42, part 1, that the following introductory statement holds true:
$\downarrow(c\cap W_{A})=\{x\in F\mid c\cap W_{A}\supseteq x\}=\{x\in F_{A}\mid c\supseteq x\}=(\downarrow c)\cap F_{A}.$
From this, the claim follows easily: If $x\in P(c\cap W_{A})$, then there is $y\in(\downarrow c)\cap F_{A}$ such that
$(\dagger)\qquad\uparrow x=\uparrow y\setminus((\downarrow c)\cap F_{A}).$
Hence, $x\in\uparrow y$. There is unique $\omega\in A$ such that $y\in(\downarrow c)\cap T_{\omega}$. As $T_{\omega}$ is a connected component of $(F,\supseteq)$ and $y\in T_{\omega}$, we have $\uparrow y\subseteq T_{\omega}$, whence $x\in T_{\omega}$ and
@@
If conversely $x\in F_{A}$ is such that there is $y\in\downarrow c$ satisfying ()($\ast$)( ∗ ), then $x\in\uparrow y$, hence $x,y$ belong to the same connected component, whence $y\in F_{A}$ and $\uparrow y\subseteq F_{A}$. We infer that ()($\dagger$)( † ) holds true, i.e., $x\in P(c\cap W_{A})$ by the introductory statement. ∎
@@
Consider the basic version $(F,\pi,{\bm{\mathsf{X}}})$ of the simple sdf over $W$ and the exogenous scenario space $(\Omega,\mathscr{A})$ as in subsection 4.3. Let $M$ be the set of maps $\Omega\to\{1,2\}$. Then the following subsets of $W$ define non-redundant and complete choices:
  • 1.
@@
Consider the variant $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ of the simple sdf over $W^{\prime}$ and the exogenous scenario space $(\Omega,\mathscr{A})$ as in subsection 4.3. Let $M$ be the set of maps $\Omega\to\{1,2\}$. Then the following subsets of $W^{\prime}$ define non-redundant and complete choices:
  • 1.
@@
If $t<u$ and $w\in c$, then any $w^{\prime}=(\omega^{\prime},f^{\prime})\in x$ satisfies $\omega^{\prime}=\omega$ and $f^{\prime}|_{[0,u)_{\mathbb{T}}}=f|_{[0,u)_{\mathbb{T}}}$, in particular $f^{\prime}|_{[0,t]}=f|_{[0,t]}$, hence $w^{\prime}\in c$. We conclude for this case that $x\subseteq c$.
@@
Consider an action path sdf $(F,\pi,{\bm{\mathsf{X}}})$ over an exogenous scenario space $(\Omega,\mathscr{A})$, and let $W=\bigcup F$. Let $t\in\mathbb{T}$ and $w\in W$. Then,
@@
If $x=x_{t}(w)$ for $w\in c$, then, by assumption AP.C1, there is $w^{\prime}\in x_{t}(w)\setminus c$. Thus, $x$ has at least two elements. In view of lemmata 4.36 and A.46 and assumption AP.W1, ()($\ast$)( ∗ ) is satisfied for $y=\{w\}$.
If, conversely, ()($\ast$)( ∗ ) is satisfied for some $y\in F$, then $x$ is a move, whence the existence of $w_{0}\in W$ and $t_{0}\in\mathbb{T}$ with $x=x_{t_{0}}(w_{0})$. Moreover, there is $w\in y$, and as $y\in\downarrow c$, we obtain $w\in c$. By ()($\ast$)( ∗ ), we know that $x\in\uparrow y$, hence also $w\in x$. Thus, $x=x_{t_{0}}(w)$. From representation ()($\ast$)( ∗ ) and lemma 4.36, we immediately get $t_{0}\leq t$. Moreover, by assumption AP.C1 on $c\in\mathscr{C}_{t}$, $x_{t}(w)$ has at least two elements. Indeed, as $w\in c$, there is $w^{\prime}\in x_{t}(w)\setminus c$. Hence, by assumption AP.W1, $x_{t}(w)\neq x_{u}(w)$ for all $u>t$. Hence, $x_{t}(w)\in\uparrow y\setminus\downarrow c=\uparrow x$, again by representation ()($\ast$)( ∗ ) and lemma 4.36. Thus, by lemma 2.21, $t=\mathfrak{t}(x_{t}(w))\leq\mathfrak{t}(x)=t_{0}$. Hence, $t_{0}=t$ and $x=x_{t}(w)$ with $w\in c$.∎
@@
Concerning completeness, let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ such that there is $\omega\in D_{\bm{\mathsf{x}}}$ with ${\bm{\mathsf{x}}}(\omega)\in P(c)$. By lemma 4.37, there is $f\in\mathbb{A}^{\mathbb{T}}$ such that $(\omega,f)\in c$ and ${\bm{\mathsf{x}}}(\omega)=x_{t}(\omega,f)$. First, we infer that $D_{t,f}\neq\emptyset$, whence $D_{\bm{\mathsf{x}}}=D_{t,f}$ and ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f)$, by proposition 2.11. Second, we infer that $x_{t}(\omega,f)\cap c\neq\emptyset$. By assumption AP.C2, for any $\omega^{\prime}\in D_{t,f}$ there is $w^{\prime}\in x_{t}(\omega^{\prime},f)\cap c$. In particular, ${\bm{\mathsf{x}}}(\omega^{\prime})=x_{t}(\omega,f)=x_{t}(w^{\prime})$. Hence, ${\bm{\mathsf{x}}}(\omega^{\prime})\in P(c)$, by lemma 4.37. We conclude that ${\bm{\mathsf{x}}}^{-1}(P(c))=D_{\bm{\mathsf{x}}}$.
@@
Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$, $t=\mathfrak{t}({\bm{\mathsf{x}}})$, and $c\in\mathscr{C}_{\bm{\mathsf{x}}}$. In particular, $c\in\mathscr{C}_{t}$, and, by lemma 4.38, $c$ is a complete and non-redundant choice. Let $\omega\in D_{\bm{\mathsf{x}}}$. Then, by definition of $\mathscr{C}_{\bm{\mathsf{x}}}$, there is $w\in{\bm{\mathsf{x}}}(\omega)\cap c$. Hence, ${\bm{\mathsf{x}}}(\omega)=x_{\mathfrak{t}(x)}(w)=x_{t}(w)$, and, by lemma 4.37, ${\bm{\mathsf{x}}}(\omega)\in P(c)$. We conclude that $c$ is available at ${\bm{\mathsf{x}}}$. ∎
@@
(First two statements):  Let $c=c(A_{<t},i,g)$. By lemma 4.38, $c$ is a non-redundant and complete choice which proves the first claim.
Next, let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ such that $c$ is available at ${\bm{\mathsf{x}}}$. By definition of ${\bm{\mathsf{X}}}$, there is $f_{0}\in\mathbb{A}^{\mathbb{T}}$ with $D_{t,f_{0}}\neq\emptyset$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$ which we fix for the entire proof.
The second claim of the theorem is also proven directly. To show this, let $\omega\in D_{\bm{\mathsf{x}}}$. Then ${\bm{\mathsf{x}}}(\omega)\in P(c)$, hence, by lemma 4.37, there is $w\in c$ with ${\bm{\mathsf{x}}}(\omega)=x_{t}(w)$. There is $f\in\mathbb{A}^{\mathbb{T}}$ such that $w=(\omega,f)$, and as $w\in c$, we have $f(t)\in A^{i,g}_{t,\omega}$. Hence, $\omega\in D$. We conclude that $D_{\bm{\mathsf{x}}}\subseteq D$.
(Helpful statements for $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$):  Let $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$. We compute the set $P(c\cap c^{\prime})$ and its preimage under ${\bm{\mathsf{x}}}$. By definition of $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, there are $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and $A_{t}^{\prime i}\in\mathscr{B}(\mathbb{A}^{i})$ such that, with $A^{\prime}_{t}=(A^{\prime}_{t,\omega})_{\omega\in\Omega}$ and $A^{\prime}_{t,\omega}=(p^{i})^{-1}(A_{t}^{\prime i})$ for all $\omega\in D_{\bm{\mathsf{x}}}$, and $A^{\prime}_{t,\omega}=\emptyset$ for all $\omega\notin D_{\bm{\mathsf{x}}}$, we have $c^{\prime}=c(A^{\prime}_{<t},A^{\prime}_{t})$, $c^{\prime}\in\mathscr{C}_{t}$, and, for all $\omega\in D_{\bm{\mathsf{x}}}$, ${\bm{\mathsf{x}}}(\omega)\cap c^{\prime}\neq\emptyset$.
@@
Thus, $f|_{[0,t)_{\mathbb{T}}}\in A_{<t}\cap A^{\prime}_{<t}$. By definition of $f$, we must have $p^{i}\circ f(t)=g(\omega)$. Hence, $(\omega,f)\in c_{0}$. By definition of $f$, we also have ${\bm{\mathsf{x}}}(\omega)=x_{t}(\omega,f)$.
@@
(First implication):  If $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurable, then, by ()($\dagger$)( † ), ${\bm{\mathsf{x}}}^{-1}(P(c\cap c^{\prime}))\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$ for all $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$.
(Second implication): Suppose conversely that ${\bm{\mathsf{x}}}^{-1}(P(c\cap c^{\prime}))\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$ for all $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, and that assumption AP.C3 is satisfied. Hence, there is $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ with $f|_{[0,t)_{\mathbb{T}}}\in A^{\prime}_{<t}$ for all $(\omega,f)\in c$, and an intersection-stable generator $\mathscr{G}(\mathbb{A}^{i})$ of $\mathscr{B}(\mathbb{A}^{i})$ such that for all $G\in\mathscr{G}(\mathbb{A}^{i})$, we have $c(A^{\prime}_{<t},A^{i,G}_{t})\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}\cup\{\emptyset\}$. Let $G\in\mathscr{G}(\mathbb{A}^{i})$ and $c^{\prime}=c(A^{\prime}_{<t},A^{i,G}_{t})$.
Assume first that $c^{\prime}=\emptyset$. Let $\omega\in D_{\bm{\mathsf{x}}}$. Then, as $c$ is available at ${\bm{\mathsf{x}}}$, ${\bm{\mathsf{x}}}(\omega)\in P(c)$. Hence, by lemma 4.37, there is $f\in\mathbb{A}^{\mathbb{T}}$ with $w=(\omega,f)\in{\bm{\mathsf{x}}}(\omega)\cap c$. In particular, using the definition of $c$ and property AP.C3(1) of $A^{\prime}_{<t}$, $p^{i}\circ f(t)=g(\omega)$ and $f|_{[0,t)_{\mathbb{T}}}\in A^{\prime}_{<t}$. As $c^{\prime}=\emptyset$, $g(\omega)\notin G$. Hence, $g|_{D_{\bm{\mathsf{x}}}}^{-1}(G)=\emptyset\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$.
Second, assume that $c^{\prime}\neq\emptyset$. Then, by definition, $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$. Represent $c^{\prime}$ as in the beginning of the “helpful statements” part above, with $A_{t}^{\prime i}=G$. Let $c_{0}$ be defined as in the “helpful statements” part.
@@
We conclude that $(g|_{D_{\bm{\mathsf{x}}}})^{-1}(G)\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$ for all $G\in\mathscr{G}(\mathbb{A}^{i})$. As $\mathscr{G}(\mathbb{A}^{i})$ is an intersection-stable generator of $\mathscr{B}(\mathbb{A}^{i})$, $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurable. ∎
@@
Regarding AP.C3, let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ be such that, with $t=\mathfrak{t}({\bm{\mathsf{x}}})$, there are $D\in\mathscr{A}$, $g\colon D\to\mathbb{A}^{i}$ and $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ satisfying $c(A_{<t},i,g)\in\mathscr{C}_{t}$ and making $c(A_{<t},i,g)$ available at ${\bm{\mathsf{x}}}$. By lemma 4.38, there is $(\omega_{0},f_{0})\in c(A_{<t},i,g)$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$ and $\omega_{0}\in D_{\bm{\mathsf{x}}}$.
@@
Let $A^{\prime}_{<t}=\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and $\mathscr{G}(\mathbb{A}^{i})=\mathscr{B}(\mathbb{A}^{i})\setminus\{\mathbb{A}^{i}\}$ which is obviously an intersection-stable generator of $\mathscr{B}(\mathbb{A}^{i})$. Clearly, $A_{<t}\subseteq A^{\prime}_{<t}$ so that property AP.C3(1) is satisfied. Let $G\in\mathscr{G}(\mathbb{A}^{i})$. If $G=\emptyset$, then $c(A^{\prime}_{<t},A_{t}^{i,G})=\emptyset$. If $G\in\mathscr{G}(\mathbb{A}^{i})\setminus\{\emptyset\}$, then the general statement above in case B) applies, hence $c(A^{\prime}_{<t},A_{t}^{i,G})\in\mathscr{C}_{t}$. To complete the proof that $c(A^{\prime}_{<t},A_{t}^{i,G})\in\mathscr{C}_{\bm{\mathsf{x}}}$, it remains to prove property (4) in the definition of $\mathscr{C}_{\bm{\mathsf{x}}}$, (3) having been proven just before and (1) and (2) being evident. For this, let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}$. Then there is $f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ such that $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f_{0}|_{[0,t)_{\mathbb{T}}}\in\mathbb{A}^{[0,t)_{\mathbb{T}}}=A^{\prime}_{<t}$ and $p^{i}\circ f^{\prime}(t)\in G$, because $G$ is non-empty and $p^{i}$ surjective. Then, $w^{\prime}=(\omega^{\prime},f^{\prime})\in\Omega\times\mathbb{A}^{\mathbb{T}}=W$, and thus $w^{\prime}\in{\bm{\mathsf{x}}}_{t}(f_{0})(\omega^{\prime})\cap c(A^{\prime}_{<t},A_{t}^{i,G})$. Property (4) therefore holds true and the proof of the statement $c(A^{\prime}_{<t},A_{t}^{i,G})\in\mathscr{C}_{\bm{\mathsf{x}}}\cup\{\emptyset\}$ is complete.
@@
(Ad timing game):  Let $t\in\mathbb{T}$. Further, let $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ be a non-empty set of component-wise decreasing paths $f_{t}$ such that $p^{i}\circ f_{t}=1_{[0,t)_{\mathbb{T}}}$, and $g\colon\Omega\to\{0,1\}=\mathbb{A}^{i}$. Let $c=c(A_{<t},i,g)$. We are going to prove that $c\in\mathscr{C}_{t}$.
@@
Regarding AP.C0, let $\omega\in\Omega$. By assumption on $A_{<t}$, there is component-wise decreasing $f\colon\mathbb{T}\to\mathbb{A}$ such that $f|_{[0,t)_{\mathbb{T}}}\in A_{<t}$ and $p^{i}\circ f(t)=g(\omega)$. Hence, $(\omega,f)\in W$, and even $(\omega,f)\in c$. Thus, $c\neq\emptyset$.

### v2

@@ A.4 Section 4
As $c\cap W_{E}=\bigcup_{\omega\in E}(c\cap W_{\omega})$ and $\{T_{\omega}\mid\omega\in\Omega\}$ is a partition of $F$, we infer from Lemma A.47, Part 1, that the following introductory statement holds true:
$\downarrow(c\cap W_{E})=\{x\in F\mid c\cap W_{E}\supseteq x\}=\{x\in F_{E}\mid c\supseteq x\}=(\downarrow c)\cap F_{E}.$
From this, the claim follows easily: If $x\in P(c\cap W_{E})$, then there is $y\in(\downarrow c)\cap F_{E}$ such that
$(\dagger)\qquad\uparrow x=\uparrow y\setminus((\downarrow c)\cap F_{E}).$
Hence, $x\in\uparrow y$. There is unique $\omega\in E$ such that $y\in(\downarrow c)\cap T_{\omega}$. As $T_{\omega}$ is a connected component of $(F,\supseteq)$ and $y\in T_{\omega}$, we have $\uparrow y\subseteq T_{\omega}$, whence $x\in T_{\omega}$ and
@@
If conversely $x\in F_{E}$ is such that there is $y\in\downarrow c$ satisfying ()($\ast$)( ∗ ), then $x\in\uparrow y$, hence $x,y$ belong to the same connected component, whence $y\in F_{E}$ and $\uparrow y\subseteq F_{E}$. We infer that ()($\dagger$)( † ) holds true, i.e. $x\in P(c\cap W_{E})$ by the introductory statement. ∎
@@
Consider the basic version $(F,\pi,{\bm{\mathsf{X}}})$ of the simple sdf on the exogenous scenario space $(\Omega,\mathscr{E})$ as in Subsection 4.3. Let $M$ be the set of maps $\Omega\to\{1,2\}$. Then the following subsets of $W$ define non-redundant and complete choices:
@@
Consider the variant $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ of the simple sdf on the exogenous scenario space $(\Omega,\mathscr{E})$ as in Subsection 4.3. Let $M$ be the set of maps $\Omega\to\{1,2\}$. Then the following subsets of $W^{\prime}$ define non-redundant and complete choices:
@@
If $t<u$ and $w\in c$, then any $w^{\prime}=(\omega^{\prime},f^{\prime})\in x$ satisfies $\omega^{\prime}=\omega$ and $f^{\prime}|_{[0,u)_{\mathbb{T}}}=f|_{[0,u)_{\mathbb{T}}}$, in particular $f^{\prime}|_{[0,t]_{\mathbb{T}}}=f|_{[0,t]_{\mathbb{T}}}$, hence $w^{\prime}\in c$. We conclude for this case that $x\subseteq c$.
@@
Consider action path sdf data $(I,\mathbb{A},\mathbb{T},W)$ on an exogenous scenario space $(\Omega,\mathscr{E})$ and the induced action path sdf $(F,\pi,{\bm{\mathsf{X}}})$. Let $t\in\mathbb{T}$ and $w\in W$. Then,
@@
If $x=x_{t}(w)$ for $w\in c$, then, by Assumption AP.C1, there is $w^{\prime}\in x_{t}(w)\setminus c$. Thus, $x$ has at least two elements. In view of Lemmata 4.39 and A.51 and Assumption AP.SDF1, ()($\ast$)( ∗ ) is satisfied for $y=\{w\}$.
If, conversely, ()($\ast$)( ∗ ) is satisfied for some $y\in\downarrow c$, then $x$ is a move, whence the existence of $w_{0}\in W$ and $t_{0}\in\mathbb{T}$ with $x=x_{t_{0}}(w_{0})$. Moreover, there is $w\in y$, and as $y\in\downarrow c$, we obtain $w\in c$. By ()($\ast$)( ∗ ), we know that $x\in\uparrow y$, hence also $w\in x$. Thus, $x=x_{t_{0}}(w)$. From representation ()($\ast$)( ∗ ) and Lemma 4.39, we immediately get $t_{0}\leq t$. Moreover, by Assumption AP.C1 on $c\in\mathscr{C}_{t}$, $x_{t}(w)$ has at least two elements. Indeed, as $w\in c$, there is $w^{\prime}\in x_{t}(w)\setminus c$. Hence, by Lemma 2.17, $\mathbb{T}_{{\bm{\mathsf{x}}}_{t}(w)}=\{t\}$. Hence, $x_{t}(w)\in\uparrow y\setminus\downarrow c=\uparrow x$, again by representation ()($\ast$)( ∗ ), the fact that $y\in\downarrow c$ and Lemma 4.39. Thus, by Lemma 2.24, $t=\mathfrak{t}(x_{t}(w))\leq\mathfrak{t}(x)=t_{0}$. Hence, $t_{0}=t$ and $x=x_{t}(w)$ with $w\in c$.∎
@@
Concerning completeness, let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ such that there is $\omega\in D_{\bm{\mathsf{x}}}$ with ${\bm{\mathsf{x}}}(\omega)\in P(c)$. By Lemma 4.40, there is $f\in\mathbb{A}^{\mathbb{T}}$ such that $(\omega,f)\in c$ and ${\bm{\mathsf{x}}}(\omega)=x_{t}(\omega,f)$. First, we infer that $D_{t,f}\neq\emptyset$, whence $D_{\bm{\mathsf{x}}}=D_{t,f}$ and ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f)$, by Proposition 2.12. Second, we infer that $x_{t}(\omega,f)\cap c\neq\emptyset$. By Assumption AP.C2, for any $\omega^{\prime}\in D_{t,f}$ there is $w^{\prime}\in x_{t}(\omega^{\prime},f)\cap c$. In particular, ${\bm{\mathsf{x}}}(\omega^{\prime})=x_{t}(\omega^{\prime},f)=x_{t}(w^{\prime})$. Hence, ${\bm{\mathsf{x}}}(\omega^{\prime})\in P(c)$, by Lemma 4.40. We conclude that ${\bm{\mathsf{x}}}^{-1}(P(c))=D_{\bm{\mathsf{x}}}$.
@@
Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$, $t=\mathfrak{t}({\bm{\mathsf{x}}})$, and $c\in\mathscr{C}_{\bm{\mathsf{x}}}^{i}$. In particular, $c\in\mathscr{C}_{t}$, and, by Lemma 4.41, $c$ is a non-redundant and complete, in particular $\tilde{\bm{\mathsf{X}}}^{i}$-complete choice. Let $\omega\in D_{\bm{\mathsf{x}}}$. Then, by definition of $\mathscr{C}_{\bm{\mathsf{x}}}^{i}$, there is $w\in{\bm{\mathsf{x}}}(\omega)\cap c$. Hence, ${\bm{\mathsf{x}}}(\omega)=x_{\mathfrak{t}({\bm{\mathsf{x}}})}(w)=x_{t}(w)$, and, by Lemma 4.40, ${\bm{\mathsf{x}}}(\omega)\in P(c)$. We conclude that $c$ is available at ${\bm{\mathsf{x}}}$. ∎
@@
Let $c=c(A_{<t},i,g)$.
(Ad 1):  By Lemma 4.41, $c$ is a non-redundant and complete, in particular $\tilde{\bm{\mathsf{X}}}^{i}$-complete choice which proves the first claim.
(Ad 2):  Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c$ is available at ${\bm{\mathsf{x}}}$. Let $\omega\in D_{\bm{\mathsf{x}}}$. Then ${\bm{\mathsf{x}}}(\omega)\in P(c)$, hence, by Lemma 4.40, there is $w\in c$ with ${\bm{\mathsf{x}}}(\omega)=x_{t}(w)$. There is $f\in\mathbb{A}^{\mathbb{T}}$ such that $w=(\omega,f)$, and as $w\in c$, we have $f(t)\in A^{i,g}_{t,\omega}$. Hence, $\omega\in D$. We conclude that $D_{\bm{\mathsf{x}}}\subseteq D$.
(Helpful statements for $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ and ${\bm{\mathsf{x}}}$ that $c$ is available at):  Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c$ is available at ${\bm{\mathsf{x}}}$. By definition of ${\bm{\mathsf{X}}}$, there is $f_{0}\in\mathbb{A}^{\mathbb{T}}$ with $D_{t,f_{0}}\neq\emptyset$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$.
Let $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$. We compute the set $P(c\cap c^{\prime})$ and its preimage under ${\bm{\mathsf{x}}}$. By definition of $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, there are $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ and $A_{t}^{\prime i}\in\mathscr{B}(\mathbb{A}^{i})$ such that, with $A^{\prime}_{t}=(A^{\prime}_{t,\omega})_{\omega\in\Omega}$ and $A^{\prime}_{t,\omega}=(p^{i})^{-1}(A_{t}^{\prime i})$ for all $\omega\in D_{\bm{\mathsf{x}}}$, and $A^{\prime}_{t,\omega}=\emptyset$ for all $\omega\notin D_{\bm{\mathsf{x}}}$, we have $c^{\prime}=c(A^{\prime}_{<t},A^{\prime}_{t})$, $c^{\prime}\in\mathscr{C}_{t}$, and, for all $\omega\in D_{\bm{\mathsf{x}}}$, ${\bm{\mathsf{x}}}(\omega)\cap c^{\prime}\neq\emptyset$.
@@
Thus, $f|_{[0,t)_{\mathbb{T}}}\in A_{<t}\cap A^{\prime}_{<t}$. By definition of $f$ and $f_{0}$, we must have $p^{i}\circ f(t)=g(\omega)$. Hence, $(\omega,f)\in c_{0}$. By definition of $f$, we also have ${\bm{\mathsf{x}}}(\omega)=x_{t}(\omega,f)$.
@@
(Ad 3):  Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c$ is available at ${\bm{\mathsf{x}}}$. If $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurable, then, by ()($\dagger$)( † ), ${\bm{\mathsf{x}}}^{-1}(P(c\cap c^{\prime}))\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$ for all $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$.
(Ad 4):  Suppose conversely that ${\bm{\mathsf{x}}}^{-1}(P(c\cap c^{\prime}))\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$ for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ that $c$ is available at and all $c^{\prime}\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, and that Assumption AP.C3 is satisfied for $(A_{<t},i,g)$ and $\mathscr{C}^{i}$.
Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c$ is available at it. By Assumption AP.C3, there is a generator $\mathscr{G}(\mathbb{A}^{i})$ of $\mathscr{B}(\mathbb{A}^{i})$, stable under non-trivial intersections, such that for all $G\in\mathscr{G}(\mathbb{A}^{i})$, we have $c(A_{<t},A^{i,G}_{t})\in\mathscr{C}^{i}_{\bm{\mathsf{x}}}$. Let $G\in\mathscr{G}(\mathbb{A}^{i})$ and $c^{\prime}=c(A_{<t},A^{i,G}_{t})$.
Represent $c^{\prime}$ as in the beginning of the “Helpful statements” part above, with $A_{t}^{\prime i}=G$ and $A^{\prime}_{<t}=A_{<t}$. Let $c_{0}$ be defined as in the “Helpful statements” part.
@@
We conclude that $(g|_{D_{\bm{\mathsf{x}}}})^{-1}(G)\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$ for all $G\in\mathscr{G}(\mathbb{A}^{i})$. Trivially, we have $(g|_{D_{\bm{\mathsf{x}}}})^{-1}(\emptyset)=\emptyset\in\mathscr{F}^{i}_{\bm{\mathsf{x}}}$. As $\mathscr{G}(\mathbb{A}^{i})\cup\{\emptyset\}$ is a generator of $\mathscr{B}(\mathbb{A}^{i})$ stable under intersections, in view of the $\pi$-$\lambda$ theorem, $g|_{D_{\bm{\mathsf{x}}}}$ is $\mathscr{F}^{i}_{\bm{\mathsf{x}}}$-measurable. ∎
@@
Regarding AP.C3, suppose that for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ contains all $c(A_{<t},A_{t})$ ranging over all $A_{t}$ satisfying ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.$k$), $k=2,3,4$. Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c(A_{<t},i,g)$ is available at ${\bm{\mathsf{x}}}$. By Lemma 4.41, there is $(\omega_{0},f_{0})\in c(A_{<t},i,g)$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$ and $\omega_{0}\in D_{\bm{\mathsf{x}}}$.
@@
Let $\mathscr{G}(\mathbb{A}^{i})=\mathscr{B}(\mathbb{A}^{i})\setminus\{\mathbb{A}^{i},\emptyset\}$ which is obviously a generator of $\mathscr{B}(\mathbb{A}^{i})$ stable under non-trivial intersections. Let $G\in\mathscr{G}(\mathbb{A}^{i})$. Then, the general statement ()($\ast$)( ∗ ) above in case B) applies, hence $c(A_{<t},A_{t}^{i,G})\in\mathscr{C}_{t}$. To complete the proof that $c(A_{<t},A_{t}^{i,G})\in\mathscr{C}_{\bm{\mathsf{x}}}^{i}$, in view of our additional assumption on $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, it remains to prove Property ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.4) for $c(A_{<t},A_{t}^{i,G})$, ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.3) having been proven just before and ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.2) being evident. For this, let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}$. Then there is $f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ such that $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f_{0}|_{[0,t)_{\mathbb{T}}}\in A_{<t}$ and $p^{i}\circ f^{\prime}(t)\in G$, because $G$ is non-empty and $p^{i}$ surjective. Then, $w^{\prime}=(\omega^{\prime},f^{\prime})\in\Omega\times\mathbb{A}^{\mathbb{T}}=W$, and thus $w^{\prime}\in{\bm{\mathsf{x}}}_{t}(f_{0})(\omega^{\prime})\cap c(A_{<t},A_{t}^{i,G})$. Property ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.4) therefore holds true for $c(A_{<t},A_{t}^{i,G})$ and the proof of the statement $c(A^{\prime}_{<t},A_{t}^{i,G})\in\mathscr{C}_{\bm{\mathsf{x}}}^{i}\cup\{\emptyset\}$ is complete.
@@
(Ad timing problem):  Let $t\in\mathbb{T}$. Further, let $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ be a non-empty set of componentwise decreasing paths $f_{t}$ such that $p^{i}\circ f_{t}=1_{[0,t)_{\mathbb{T}}}$, and $g\colon\Omega\to\{0,1\}=\mathbb{A}^{i}$. Let $c=c(A_{<t},i,g)$. We are going to prove that $c\in\mathscr{C}_{t}$.
@@
Regarding AP.C0, let $\omega\in\Omega$. By assumption on $A_{<t}$, there is componentwise decreasing $f\colon\mathbb{T}\to\mathbb{A}$ such that $f|_{[0,t)_{\mathbb{T}}}\in A_{<t}$ and $p^{i}\circ f(t)=g(\omega)$. Hence, $(\omega,f)\in W$, and even $(\omega,f)\in c$. Thus, $c\neq\emptyset$.